OpenStudy (anonymous):

find the solution of this sytsem of equations8x-4y=-32 & -3x+4y=-42

OpenStudy (anonymous):

Add both equations together, this will eliminate the y variable and you'll just have to solve for x. Then substitute the value of x back into one of the equations and solve for y.

OpenStudy (anonymous):

makes ni sence

OpenStudy (anonymous):

First step. Add both equations together: 8x -4y = -32 +(-3x+4y=-42) ----------------- What do you get?

OpenStudy (anonymous):

like add what and what together -32 and -42

OpenStudy (anonymous):

All of them: 8x-3x=how many x's -4y+4y=how many y's -32-42= what number

OpenStudy (anonymous):

X=5 Y=0 AND NUMBERS=10

OpenStudy (anonymous):

With the new values your form a new equation. 5x=10 Next step: Solve for x

OpenStudy (anonymous):

would it be 5?

OpenStudy (anonymous):

No. According to the equation it's a number that multiplied times 5 will give you 10.

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

Right. Now next step. Substitute that value in place of the x's in either one of the two original equations. 8x-4y=-32 8(2)-4y=-32 Do the operation, then solve for y.